# Shaping Realities: Enhancing 3D Generative AI with Fabrication Constraints

### v1

@@ 3. Encoding Physical Functionality in Generative 3D Models
A critical aspect in the evolution of 3D modeling, especially within the context of generative AI, is encoding physical functionality into the generated designs. As AI-based tools are increasingly used for creating 3D models, a significant challenge emerges: ensuring that these models are not just visually appealing but also functionally viable when fabricated. The process involves integrating complex physical principles and material properties into the generative algorithms, which goes beyond the traditional focus on aesthetic design.
The integration of physical functionality is especially crucial when considering the end use of these models. Generative AI systems, such as TextMesh (Tsalicoglou et al., 2024) and DreamGaussian (Tang et al., 2023) are typically adept at creating detailed and intricate designs. However, without the encoding of physical functionality, these designs may lack practical applicability, especially when subjected to specific loading conditions including static and dynamic forces, pressure, and other environmental factors.
Addressing this challenge involves several key considerations:
@@
The goal of encoding physical functionality in generative 3D models is to bridge the gap between aesthetically driven design and practical, real-world applicability. This approach not only enhances the usability of the generated models but also opens up new possibilities for innovation in various fields where functionality is as critical as form. By addressing these challenges, we can pave the way for generative AI to play a pivotal role in the future of 3D modeling and manufacturing, where functionality is encoded at the very core of the design process.

### v2

@@ 3. Encoding Physical Functionality in Generative 3D Models
A critical aspect in the evolution of 3D modeling, especially within the context of generative AI, is encoding physical functionality into the generated designs. As AI-based tools are increasingly used for creating 3D models, a significant challenge emerges: ensuring that these models are not just visually appealing but also functionally viable when fabricated. The process involves integrating complex physical principles and material properties into the generative algorithms, which goes beyond the traditional focus on aesthetic design. Generative AI systems (Tsalicoglou et al., 2024; Tang et al., 2023) are typically adept at creating detailed and intricate designs. However, without the encoding of physical functionality, these designs may lack practical applicability, especially when subjected to specific loading conditions including static and dynamic forces, pressure, and other environmental factors. Addressing this challenge involves several key considerations:
@@
The goal of encoding physical functionality in generative 3D models is to bridge the gap between aesthetically driven design and practical, real-world applicability. This approach not only enhances the usability of the generated models but also opens up new possibilities for innovation in various fields where functionality is at the core of the design process.
@@ 4. Conclusion
5. Acknowledgments
We thank the MIT-Google Program for Computing Innovation for their generous support. Furthermore, we thank Douglas Eck and Yuanzhen Li from Google for their support to this research.